Status Report

ISS On-Orbit Status 28 June 2002

By SpaceRef Editor
June 28, 2002
Filed under , ,

All ISS systems continue to function nominally,
except as noted previously or below.

The test of the Service Module’s Kurs-P subset 2 was successfully
conducted, concluding at about 3:20am EDT.  It used the Progress
8P, trailing the ISS at about 30 km distance, as active vehicle.
MCC-Moscow collected a large amount of data, and its analysis is
underway.  Early reports are that the Kurs equipment on the SM
surprisingly appeared to "behave much better" than during
the docking of Progress 7P on 3/24/02.

After the test, the DV3 burn of Progress 8P was executed without a
hitch at about 5:55am EDT, and the cargo ship is now on its final
approach trajectory, closing in from below and behind the ISS.
 [The station continues in XPOP attitude, with the 2B and 4B
solar wings feathered to 219 deg and 141 deg, respectively, in
preparation for Progress’ arrival during the night.  Start of
the terminal approach (space-to-space comm established) is at 1:40am,
at a range of 8 km. With three additional thruster burns, 8P will
close the distance, until docking occurs at 2:25 am EDT (9:25am DMT).
 The event will be carried on NASA TV.]

Valery Korzun and Peggy Whitson configured the television
equipment for downlinking the docking event to the ground and also
for videotaping it on board.  The downlink was then tested at
8:50am EDT. [The config provides for the routing of video signals
from the RS (Russian segment) to the USOS (US segment) and subsequent
transmittal to MCC-H/MCC-M) on Ku-band via TDRS.]

FE-2 Sergei Treschev performed the first monthly EVARM (EVA
radiation monitoring) background radiation badge reading, to measure
radiation levels on station.  [After setting up the EVARM
reader and applying power from the UOP (utility outlet panel), a
15-minute delay in the schedule was required for warm-up of the
reader. All 12 dosimeter badges were read, and the data were
downloaded to the HRF (human research facility) PC. Treschev then
connected the PC to the HRF Rack, which POC (Payload Operations
Center, Huntsville) had powered up and configured in preparation for
downlinking of the EVARM data and Wednesday’s PuFF (pulmonary
function in flight) files.  After the downlink Sergei
deactivated the rack.]

For FE-1 Peggy Whitson, today was the first ADVASC (advanced
astroculture) nutrient exchange.  In this procedure, she removed
the now diluted nutrient fluid from the system and added fresh fluid
with the full concentration of nutrients, with supportive ground

Peggy later downloaded the crewmembers’ exercise files from the TVIS
(treadmill), RED (resistive exerciser), and CEVIS (cycle ergometer)
to the MEC (medical equipment computer) for subsequent transmittal to
the ground.

Sergei Treschev started absorbent bed #1 of the BMP micropurification
unit on the regular 24-hr. regeneration bake-out, with filter channel
#2 continuing in absorption mode.

While Peggy handled today’s autonomous payloads status check in the
Lab, CDR Valery Korzun performed the daily maintenance routine on the
SOSh environment control and life support system and prepared today’s
IMS (inventory management system) "delta" file.

All crewmembers performed their daily physical exercise program.

Yesterday, the crew installed the new MSA (mass spectrometer
assembly) and VGA (verification gas assembly) in the MCA (major
constituent analyzer) for air sample analysis.  The task had
been on the "job jar" task list.  For the MCA to
return to full functionality, a second task of mating connectors and
electronics at the rear of the rack will be scheduled in about two

The VOA (volatile organics analyzer) yesterday had a "phantom
reboot", i.e., it booted up without ground commanding or crew
action and went to standby mode.  After the data dumps were
completed, the VOA did not respond to ground commands to terminate
the data dump. MCC-H believes that the instrument is locked up.
 The received data are good and are being used for VOA

Yesterday at about 3:00 am EDT, CMG2 (control moment gyroscope #2)
experienced a loss of communications, apparently a repeat of the
event that occurred last weekend.  No other offnominal parameter
behavior was seen.  Comm was soon restored, and specialists are
investigating this latest event.  The three functioning CMGs are
continuing to work fine.

Repair of the failed CDRA (carbon dioxide removal assembly) in the
Lab is now scheduled for 7/15-16.

The condensate tank in the Lab is filling up again, currently with a
little less than 80 lbs of water.  A condensate dump into CWC
(collapsible water container) bags is planned for next week.

The two Russian Orlan-EVAs for Increment 5 are now scheduled for 8/14
and 8/20 (there is also a backup plan).  This will be acceptable
to MCC-H if, as it becomes increasingly likely, the launch of
STS-112/9A is delayed from 8/22 due to the investigation of the
cracks discovered in Orbiter LH2 pipe bellows flow liners.

Moscow is currently in the process of certifying the SM for
performing station reboosts with its thrusters.  The last (and
only) time SM propulsion was used for this purpose was during the
early SM-alone phase of the assembly sequence.  Specialists of
the joint "Team Zero" are also studying the use of SM
thrusters for performing orbital debris avoidance maneuvers,
currently the function of Progress ships.  They are also looking
into the possibility of leaving one Progress at the ISS at all times,
i.e., with one Progress docked at the SM aft port, the other at the
DC-1 docking module.  The first Progress under consideration
would be M1-10 (11P) in 2003.  At this time, these "Team
Zero" activities are only conceptual studies.

Investigation is underway of the persistent discrepancy between the
on-board CO2 readings of the U.S. CDMK (carbon dioxide monitoring
kit) and the Russian CO2 telemetry data from the SM.

Today’s targets for CEO (crew earth observation) were European
(with stagnant high pressure in place over the central
Med, this was an excellent pass for documenting aerosol buildup over
Italy. Of interest, as ISS tracked down the W coast of the peninsula:
oblique and limb views to the right of track [NE] to document the
thickness and extent of the smog there)
, Lake Nasser, Toshka
Lakes; Egypt
(after years of high volume flow for the Nile
River, concern is growing that a period of lower volume may now be
starting and will further strain Egypt’s water supply. On this pass
the crew was to use the ESC [electronic still camera] and the long
lenses to map details of the river below the Aswan Dam, noting the
water color and base lining the current land use patterns there)
Missionary Ridge, CO Fires (Dynamic Event Site: The latest
large fire in Colorado is located in the San Juan National Forest in
southwestern part of the state. Of interest: looking to the left of
track for views of this 67,000+acre fire. Additional fires to the
northwest in eastern Utah may also have been visible)
, Canary
Cloud Vortices
(Dynamic Event Site:  A large field of
stratus clouds was advancing southwestward towards the islands. Crew
was to look to the right of track for large wake cloud vortices
forming south and west of the islands)
, Canadian Rocky
(fair summer weather holds over W central Canada. Of
interest on this pass: oblique, context views to the left of track
[N], particularly noting the remaining patches of snow and ice. using
the long lenses of the ESC here)

CEO images can be viewed at the website

ISS Orbit (as of this morning, 7:41pm EDT [=
Mean altitude — 392.5 km
Apogee — 397.4 km
Perigee — 387.5 km
Period — 92.4 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0007315
Orbits per 24-hr. day — 15.58
Altitude decrease — 200 m (mean) in last 24 hours
Solar Beta Angle —  -27.0 deg (magnitude decreasing)
Revolutions since FGB/Zarya launch (Nov. ’98) — 20579
Current Flight Attitude — XPOP (x-axis perpendicular to orbit
plane = “sun-fixed” [yaw: -1.2 deg, pitch: -7.4 deg., roll: 0

For more on ISS orbit and naked-eye visibility dates/times, see

SpaceRef staff editor.